OpenStudy (dls):

\[\lim_{x \rightarrow \-3} \frac{x^5+243}{x^4-81}\]

OpenStudy (dls):

\[\lim_{x \rightarrow -3 } \frac{x^5+243}{x^4-81}\]

OpenStudy (ajprincess):

\(x^5+243=x^5+3^5\) \(=(x+3)(x^4-3x^3+9x^2-27x+81)\) \(x^4-81=x^4-3^4\) \(=(x^2-3^2)(x^2+3^2)\) \(=(x-3)(x+3)(x^2+9)\) \[\lim_{x \rightarrow -3 } \frac{x^5+243}{x^4-81}\] \[=\lim_{x \rightarrow -3 } \frac{(x+3)(x^4-3x^3+9x^2-27x+81)}{(x-3)(x+3)(x^2+9)}\] \[=\lim_{x \rightarrow -3 }\frac{(x^4-3x^3+9x^2-27x+81)}{(x-3)(x^2+9)}\] Substitute x=-3 in it to find the value

OpenStudy (dls):

You did a complex and lengthy way though

OpenStudy (dls):

Anyone with a shorter one is still open

OpenStudy (anonymous):

Use L'Hopital.

OpenStudy (anonymous):

5x/4

OpenStudy (dls):

5x^4/4x^3

OpenStudy (anonymous):

= 15/4

OpenStudy (dls):

where did the powers go?

OpenStudy (anonymous):